3-(3x/2)=8-4x
Solve for x

Answer by Cromlix(4381)

3-(3x/2)=8-4x
3 - 3x/2 = 8 - 4x
Denominator = 2
6 - 3x = 16 - 8x
-3x + 8x = 16 - 6
5x = 10
x = 2.
Hope this helps.
:-)